Diagonalize the matrix

10 −2 −5

−2 2 3

−5 3 5


1
Expert's answer
2021-10-11T02:58:32-0400
"A=\begin{pmatrix}
 10 & -2 & -5 \\
 -2 & 2 & 3 \\
 -5 & 3 & 5 \\
\end{pmatrix}"

Find the eigen values


"A-\lambda I=\begin{pmatrix}
 10-\lambda & -2 & -5 \\
 -2 & 2-\lambda & 3 \\
 -5 & 3 & 5-\lambda \\
\end{pmatrix}"

"\det(A-\lambda I)=\begin{vmatrix}
 10-\lambda & -2 & -5 \\
 -2 & 2-\lambda & 3 \\
 -5 & 3 & 5-\lambda \\
\end{vmatrix}"

"=(10-\lambda\begin{vmatrix}
 2-\lambda & 3 \\
 3 & 5-\lambda
\end{vmatrix}+2\begin{vmatrix}
 -2 & 3 \\
 -5 & 5-\lambda
\end{vmatrix}"

"-5\begin{vmatrix}
 -2 & 2-\lambda \\
 -5 & 3
\end{vmatrix}=(10-\lambda)(10-7\lambda+\lambda^2-9)"

"+2(-10+2\lambda+15)-5(-6+10-5\lambda)"

"=10-70\lambda+10\lambda^2-\lambda+7\lambda^2-\lambda^3+10+4\lambda"

"-20+25\lambda=-\lambda^3+17\lambda^2-42\lambda"

"=-\lambda(\lambda-3)(\lambda-14)"

"\det(A-\lambda I)=0=>-\lambda(\lambda-3)(\lambda-14)=0"



The roots are "\lambda_1=14, \lambda_2=3, \lambda_3=0."

These are eigenvalues.

Find the eigenvectors

"\lambda
=14"


"A-\lambda I=\begin{pmatrix}
 10-14 & -2 & -5 \\
 -2 & 2-14 & 3 \\
 -5 & 3 & 5-14 \\
\end{pmatrix}"

"=\begin{pmatrix}
 -4 & -2 & -5 \\
 -2 & -12 & 3 \\
 -5 & 3 & -9 \\
\end{pmatrix}"

"R_2=R_2-R_1\/2"


"\begin{pmatrix}
 -4 & -2 & -5 \\
 0 & -11 & 11/2 \\
 -5 & 3 & -9 \\
\end{pmatrix}"

"R_3=R_3-5R_1\/4"


"\begin{pmatrix}
 -4 & -2 & -5 \\
 0 & -11 & 11/2 \\
 0 & 11/2 & -11/4 \\
\end{pmatrix}"

"R_3=R_3+R_2\/2"


"\begin{pmatrix}
 -4 & -2 & -5 \\
 0 & -11 & 11/2 \\
 0 & 0 & 0 \\
\end{pmatrix}"

"R_2=R_2\/(-11)"


"\begin{pmatrix}
 -4 & -2 & -5 \\
 0 & 1 & -1/2 \\
 0 & 0 & 0 \\
\end{pmatrix}"

"R_1=R_1+2R_2"


"\begin{pmatrix}
 -4 & 0 & -6 \\
 0 & 1 & -1/2 \\
 0 & 0 & 0 \\
\end{pmatrix}"

"R_1=R_1\/(-4)"


"\begin{pmatrix}
 1 & 0 & 3/2 \\
 0 & 1 & -1/2 \\
 0 & 0 & 0 \\
\end{pmatrix}"

If we take "v_3=t," then "v_1=-\dfrac{3}{2}t, v_2=\dfrac{1}{2}t."

The eigenvector is "v=\begin{pmatrix}
 - 3/2 \\
 1/2 \\
1
\end{pmatrix}"


"\lambda
=3"

"A-\lambda I=\begin{pmatrix}
 10-3 & -2 & -5 \\
 -2 & 2-3 & 3 \\
 -5 & 3 & 5-3 \\
\end{pmatrix}"

"=\begin{pmatrix}
 7 & -2 & -5 \\
 -2 & -1 & 3 \\
 -5 & 3 & 2 \\
\end{pmatrix}"

"R_2=R_2+2R_1\/7"

"\begin{pmatrix}
 7 & -2 & -5 \\
 0 & -11/7 & 11/7 \\
 -5 & 3 & 2 \\
\end{pmatrix}"

"R_3=R_3+5R_1\/7"

"\begin{pmatrix}
 7 & -2 & -5 \\
 0 & -11/7 & 11/7 \\
 0 & 11/7 & -11/7 \\
\end{pmatrix}"

"R_3=R_3+R_2"

"\begin{pmatrix}
 7 & -2 & -5 \\
 0 & -11/7 & 11/7 \\
 0 & 0 & 0 \\
\end{pmatrix}"

"R_2=-7R_2\/11"

"\begin{pmatrix}
 7 & -2 & -5 \\
 0 & 1 &-1 \\
 0 & 0 & 0 \\
\end{pmatrix}"

"R_1=R_1+2R_2"

"\begin{pmatrix}
 7 & 0 & -7 \\
 0 & 1 &-1 \\
 0 & 0 & 0 \\
\end{pmatrix}"

"R_1=R_1\/7"

"\begin{pmatrix}
 1 & 0 & -1 \\
 0 & 1 &-1 \\
 0 & 0 & 0 \\
\end{pmatrix}"

If we take "u_3=t," then "u_1=t, u_2=t."

The eigenvector is "u=\begin{pmatrix}
 1 \\
 1 \\
1
\end{pmatrix}"



"\lambda
=0"

"A-\lambda I=\begin{pmatrix}
 10-0 & -2 & -5 \\
 -2 & 2-0 & 3 \\
 -5 & 3 & 5-0 \\
\end{pmatrix}"

"=\begin{pmatrix}
 10 & -2 & -5 \\
 -2 & 2 & 3 \\
 -5 & 3 & 5 \\
\end{pmatrix}"

"R_2=R_2+R_1\/5"

"\begin{pmatrix}
 10 & -2 & -5 \\
 0 & 8/5 &2 \\
 -5 & 3 & 5 \\
\end{pmatrix}"

"R_3=R_3+R_1\/2"

"\begin{pmatrix}
 10 & -2 & -5 \\
 0 & 8/5 &2 \\
 0 & 2 & 5/2 \\
\end{pmatrix}"

"R_3=R_3-5R_2\/4"

"\begin{pmatrix}
 10 & -2 & -5 \\
 0 & 8/5 &2 \\
 0 & 0 & 0 \\
\end{pmatrix}"

"R_2=5R_2\/8"

"\begin{pmatrix}
 10 & -2 & -5 \\
 0 & 1 & 5/4 \\
 0 & 0 & 0 \\
\end{pmatrix}"

"R_1=R_1+2R_2"

"\begin{pmatrix}
 10 & 0 & -5/2 \\
 0 & 1 & 5/4 \\
 0 & 0 & 0 \\
\end{pmatrix}"

"R_1=R_1\/10"

"\begin{pmatrix}
 1 & 0 & -1/4 \\
 0 & 1 & 5/4 \\
 0 & 0 & 0 \\
\end{pmatrix}"

If we take "w_3=t," then "w_1=\dfrac{1}{4}t, w_2=-\dfrac{5}{4}t."

The eigenvector is "u=\begin{pmatrix}
 1/4\\
 -5/4\\
1
\end{pmatrix}"

Form the matrix "P"


"P=\begin{pmatrix}
 -3/2 & 1 & 1/4 \\
 1/2 & 1 & -5/4 \\
 1 & 1 & 1
\end{pmatrix}"

Form the diagonal matrix "D"


"D=\begin{pmatrix}
 14 & 0 & 0 \\
 0 & 3 & 0 \\
 0 & 0 & 0
\end{pmatrix}"

The matrices "P" and "D" are such that the initial matrix


"A=\begin{pmatrix}
 10 & -2 & -5 \\
 -2 & 2 & 3 \\
 -5 & 3 & 5 \\
\end{pmatrix}=PDP^{-1}"

"P=\begin{pmatrix}
 -3/2 & 1 & 1/4 \\
 1/2 & 1 & -5/4 \\
 1 & 1 & 1
\end{pmatrix}"

"D=\begin{pmatrix}
 14 & 0 & 0 \\
 0 & 3 & 0 \\
 0 & 0 & 0
\end{pmatrix}"